OpenStudy (anonymous):

find all triples (a,b,c) with a,b, and c are positive integer such that (1 + 1/a)(1 + 1/b)(1 + 1/c) = 2

OpenStudy (anonymous):

Solve for: Let's solve for a. (1+ 1 a )(1+ 1 b )(1+ 1 c )=2 Step 1: Multiply both sides by abc. abc+ab+ac+bc+a+b+c+1=2abc Step 2: Add -2abc to both sides. abc+ab+ac+bc+a+b+c+1+−2abc=2abc+−2abc −abc+ab+ac+bc+a+b+c+1=0 Step 3: Add -bc to both sides. −abc+ab+ac+bc+a+b+c+1+−bc=0+−bc −abc+ab+ac+a+b+c+1=−bc Step 4: Add -b to both sides. −abc+ab+ac+a+b+c+1+−b=−bc+−b −abc+ab+ac+a+c+1=−bc−b Step 5: Add -c to both sides. −abc+ab+ac+a+c+1+−c=−bc−b+−c −abc+ab+ac+a+1=−bc−b−c Step 6: Add -1 to both sides. −abc+ab+ac+a+1+−1=−bc−b−c+−1 −abc+ab+ac+a=−bc−b−c−1 Step 7: Factor out variable a. a(−bc+b+c+1)=−bc−b−c−1 Step 8: Divide both sides by -bc+b+c+1. a(−bc+b+c+1) −bc+b+c+1 = −bc−b−c−1 −bc+b+c+1 a= −bc−b−c−1 −bc+b+c+1 Answer: a= −bc−b−c−1 −bc+b+c+1

OpenStudy (anonymous):

@rational i need help :D

OpenStudy (anonymous):

\[(1 + \frac{ 1 }{a })(1 + \frac{ 1 }{b })(1 + \frac{ 1 }{c }) = 2\]

OpenStudy (anonymous):

\[(\frac{ a+1 }{ a })(\frac{ b+1 }{ b })(\frac{ c+1 }{ c }) = 2\]

OpenStudy (anonymous):

(a + 1)(b + 1)(c + 1) = 2abc

OpenStudy (anonymous):

Solve for: Let's solve for a. (a+1)(b+1)(c+1)=2abc Step 1: Add -2abc to both sides. abc+ab+ac+bc+a+b+c+1+−2abc=2abc+−2abc −abc+ab+ac+bc+a+b+c+1=0 Step 2: Add -1 to both sides. −abc+ab+ac+bc+a+b+c+1+−1=0+−1 −abc+ab+ac+bc+a+b+c=−1 Step 3: Add -c to both sides. −abc+ab+ac+bc+a+b+c+−c=−1+−c −abc+ab+ac+bc+a+b=−c−1 Step 4: Add -b to both sides. −abc+ab+ac+bc+a+b+−b=−c−1+−b −abc+ab+ac+bc+a=−b−c−1 Step 5: Add -bc to both sides. −abc+ab+ac+bc+a+−bc=−b−c−1+−bc −abc+ab+ac+a=−bc−b−c−1 Step 6: Factor out variable a. a(−bc+b+c+1)=−bc−b−c−1 Step 7: Divide both sides by -bc+b+c+1. a(−bc+b+c+1) −bc+b+c+1 = −bc−b−c−1 −bc+b+c+1 a= −bc−b−c−1 −bc+b+c+1 Answer: a= −bc−b−c−1 −bc+b+c+1

OpenStudy (dan815):